• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Snyder

    Log entry time 00:23:54 on June 9,2004

    Entry number 125012

    keyword=Swing Shift summary

    Pulsed beam was delivered starting around 16:00. After MCC was done tuning and verified no problems of 100 uA delivery to the hall, we started with 5uA unrastered beam with RF FB and FFB ON.

    We don't think the target camera is working. We have changed from empty to BeO to Carbon targets and all monitors look the same.

    A HARP scan was done to verify beam size at the target. The BeO viewer was supposed to be used to do this but it was not working. The Hall Target OTR gave us results for spot size that we didn't believe. Thus, we did a HARP scan.

    At this time, Rob is starting optics studies with the Carbon target in with 5uA of unrastered beam.